# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 2.2.3 Linguistic Coverage
The linguistic coverage of fairness datasets has profound implications for fairness evaluation beyond mere language representation. First, language structure directly influences how biases manifest—grammatical gender in languages like Spanish or French creates different fairness challenges than those in non-gendered languages like English. Second, cultural contexts embedded in different languages shape distinct stereotype patterns, requiring culturally-sensitive evaluation approaches. Finally, models trained predominantly on high-resource languages often exhibit substantial performance disparities when evaluating fairness in low-resource languages, potentially amplifying existing inequities across linguistic communities.
@@ 2.2.4 Bias Typology
Fairness evaluations in LMs require precise characterization of the types of biases being assessed. Previous research has often employed inadequate or incomplete definitions of bias [61], frequently focusing on narrow aspects such as specific gender stereotypes [92] or gender-occupation associations [79, 62]. We propose a comprehensive bias typology that encompasses both demographic characteristics and dataset construction factors, providing a more systematic framework for fairness evaluation.
@@
i) Demographic Characteristic Biases. These biases relate to protected attributes and social identities, largely based on categories defined by the US Equal Employment Opportunities Commissionhttps://www.eeoc.gov/prohibited-employment-policiespractices. They include gender bias (WinoBias [62], WinoGender [79]), racial and ethnic bias (CrowS-Pairs [56]), age bias (Bold [77]), religion bias, disability bias, socioeconomic bias, sexual orientation bias (WinoQueer [69]), nationality and cultural bias, physical appearance bias, and intersectional bias. Each category represents a distinct dimension along which LMs may exhibit unfair associations or disparate performance. By categorizing datasets according to these demographic dimensions, researchers can target specific bias concerns in their evaluations.
@@
ii) Dataset Construction Biases. These biases relate to how fairness datasets themselves are constructed and may influence evaluation results. They include selection bias (non-representative sampling), annotation bias (subjective judgments from human annotators), stereotype reinforcement (framing that inadvertently reinforces stereotypes), cultural context bias (failure to account for cultural differences), temporal bias (outdated social norms), and spurious correlations (misleading statistical patterns). Awareness of these methodological biases is crucial for interpreting fairness evaluation results, as they can significantly influence conclusions about model performance across demographic groups. This comprehensive typology enables more systematic and targeted fairness evaluations. By clearly differentiating between demographic characteristic biases and dataset construction biases, researchers can better identify the specific mechanisms through which unfairness manifests in LMs. Additionally, this framework facilitates more precise dataset selection, allowing researchers to choose evaluation resources that align with their specific fairness concerns and application contexts.
2.2.5 Accessibility
@@
Datasets can generally be categorized based on their accessibility into two groups: (1) publicly available datasets and (2) restricted-access datasets, though the majority fall under the public category.
@@
Publicly available datasets are accessible to researchers and developers without major barriers. These datasets are often hosted on platforms like Hugging Face Datasetshttps://huggingface.co or GitHub repositorieshttps://github.com and come with documentation and licensing terms that encourage open use. They are essential for promoting transparency, collaboration, and replicability in fairness research, enabling broad participation in the evaluation and improvement of LMs.
@@ 2.3 Bias Definitions and Measurement Framework
Building on the dataset taxonomy introduced in Section 2.2—defined along the dimensions of Structure, Source, Linguistic Coverage, Bias Typology, and Accessibility—we adopt a common notation that abstracts away these surface characteristics. Concretely, every dataset is represented as a finite multiset of linguistic items $\mathcal{D}=\{x_{1},\dots,x_{N}\}$, where each instance $x_{i}$ may carry a protected attribute $A\!\in\!\mathcal{A}$ (e.g., gender, race), an optional gold label $y\!\in\!\mathcal{Y}$, or a continuous score $s\!\in\![0,1]$. Let $P_{\mathcal{P}}$ denote the target population distribution over attributes against which representativeness is judged. Within this unified formal framework, equally applicable to both minimally paired and naturally occurring text, we now define four dataset-level bias phenomena and their principled estimators.
@@
We define representativeness bias using the Kullback–Leibler (KL) divergence between the dataset distribution $P_{\mathcal{D}}(A)$ and population distribution $P_{\mathcal{P}}(A)$:
$B_{\mathrm{rep}}=D_{\mathrm{KL}}(P_{\mathcal{D}}(A)\;\|\;P_{\mathcal{P}}(A))=\sum_{a\in\mathcal{A}}P_{\mathcal{D}}(a)\log\frac{P_{\mathcal{D}}(a)}{P_{\mathcal{P}}(a)}.$ (1)
For instance, the BBQ dataset underrepresents Indigenous identities ($<2\%$) relative to U.S. census estimates ($\sim$5%), yielding a non-zero $B_{\mathrm{rep}}$.
Annotation Bias. Annotation bias occurs when human annotators or labeling algorithms assign different labels to semantically equivalent content depending on the demographic group referenced. Let $g_{\theta}(x)$ be the labeling function (human or automated) and $A$ the protected attribute.
We estimate annotation bias as the maximum expected difference in label values between demographic groups:
$B_{\text{ann}}=\max_{a_{1},a_{2}\in\mathcal{A}}\left|\mathbb{E}[g_{\theta}(x)\mid A=a_{1}]-\mathbb{E}[g_{\theta}(x)\mid A=a_{2}]\right|.$ (2)
@@
$\hat{B}_{\mathrm{ann}}=\max_{a_{1},a_{2}\in\mathcal{A}}\left|\mathbb{E}_{x\sim\mathcal{D}_{a_{1}}}[y]-\mathbb{E}_{x\sim\mathcal{D}_{a_{2}}}[y]\right|.$ (3)
This formulation captures systematic differences in labeling outcomes across identity groups. As an example, in the widely-used hate-speech corpus of Davidson et al. [93], crowdworkers labeled tweets written in African-American English (AAE) as “offensive” nearly twice as often as Standard American English (SAE) tweets with parallel semantics. As those labels become part of the dataset, any model trained or evaluated on the corpus inherits this annotation bias.
@@
$\mathrm{PMI}(w_{g},w_{t})\;=\;\log\frac{P(w_{g},w_{t})}{P(w_{g})\,P(w_{t})},\qquad w_{g}\in W_{G},\;w_{t}\in W_{T}.$ (4)
Corpus-level association. Mutual Information (MI) averages PMI over all group–trait pairs, yielding a single scalar that summarizes stereotype leakage for the whole dataset:
$\mathrm{MI}(W_{G};W_{T})\;=\;\sum_{w_{g}\in W_{G}}\;\sum_{w_{t}\in W_{T}}P(w_{g},w_{t})\,\log\frac{P(w_{g},w_{t})}{P(w_{g})\,P(w_{t})}.$ (5)
@@
Differential Metric Bias. Differential metric bias arises when scoring functions (e.g., for toxicity, sentiment, regard) yield significantly different outputs across protected groups, even for semantically comparable inputs. Such disparities may reflect flaws in the metric itself or biased correlations in the dataset.
Let $f:\mathcal{X}\rightarrow[0,1]$ be a scoring function applied to instances $x$. We define bias as the score expectation gap between groups:
$\Delta_{f}(a_{1},a_{2})=\mathbb{E}_{x\sim\mathcal{D}_{a_{1}}}[f(x)]-\mathbb{E}_{x\sim\mathcal{D}_{a_{2}}}[f(x)].$ (8)
We test the null hypothesis $\Delta_{f}(a_{1},a_{2})=0$ using parametric ($t$-test) or non-parametric (Mann–Whitney $U$) methods, and compute effect size via Cohen’s $d$:
$d=\frac{\Delta_{f}(a_{1},a_{2})}{\sigma_{\mathrm{pooled}}},\quad\sigma_{\mathrm{pooled}}=\sqrt{\frac{\sigma^{2}_{a_{1}}+\sigma^{2}_{a_{2}}}{2}}.$ (9)
For example, when scoring BOLD prompts, regard scores for sentences referencing “Muslim” individuals were systematically lower than those referencing “Christian” individuals, despite the controlled structure indicating potential metric bias.
@@
3 Counterfactual Datasets
Counterfactual datasets represent a significant category of resources designed to evaluate fairness in LMs. These datasets operate by measuring how model responses or likelihoods change when specific identity-related attributes or stereotypical associations are altered within otherwise identical or minimally differing contexts. This controlled approach allows researchers to isolate and quantify biases related to particular demographic characteristics, such as gender, race, or occupation, which ideally should not influence model behavior in many tasks. Counterfactual datasets are particularly valuable for probing biases in applications like coreference resolution, detecting harmful stereotype associations, and assessing fairness in sentiment analysis or text classification. This section provides a detailed analysis of prominent counterfactual datasets, beginning with a comparative overview based on the taxonomy established in Section 2.2, followed by in-depth examinations of individual datasets categorized by their primary evaluation focus.

### v2

@@ 2.1.2.2  Linguistic Coverage
The linguistic coverage of fairness datasets has profound implications for fairness evaluation beyond mere language representation. Second, cultural contexts embedded in different languages shape distinct stereotype patterns, requiring culturally-sensitive evaluation approaches. Finally, models trained predominantly on high-resource languages often exhibit substantial performance disparities when evaluating fairness in low-resource languages, potentially amplifying existing inequities across linguistic communities.
@@ 2.1.2.3  Bias Typology
Beyond source and linguistic coverage, each dataset targets a specific bias typology: gender, race, religion, socio‑economic status, intersectional categories, or toxicity and sentiment imbalances. Fairness evaluations in LMs require precise characterization of the types of biases being assessed. Previous research has often employed inadequate or incomplete definitions of bias [23], frequently focusing on narrow aspects such as specific gender stereotypes [50] or gender-occupation associations [51, 24]. We propose a comprehensive bias typology that encompasses both demographic characteristics and dataset construction factors, providing a more systematic framework for fairness evaluation.
@@
i) Demographic Characteristic Biases. These biases relate to protected attributes and social identities, largely based on categories defined by the US Equal Employment Opportunities Commission [^2] https://www.eeoc.gov/prohibited-employment-policiespractices. They include gender bias (WinoBias [24], WinoGender [51]), racial and ethnic bias (CrowS-Pairs [18]), age bias (Bold [36]), religion bias, disability bias, socioeconomic bias, sexual orientation bias (WinoQueer [35]), nationality and cultural bias, physical appearance bias, and intersectional bias. Each category represents a distinct dimension along which LMs may exhibit unfair associations or disparate performance. By categorizing datasets according to these demographic dimensions, researchers can target specific bias concerns in their evaluations.
@@
ii) Dataset Construction Biases. These biases relate to how fairness datasets themselves are constructed and may influence evaluation results. They include representative bias (non-representative sampling), annotation bias (subjective judgments from human annotators or external metrics), and stereotype leakage (framing that inadvertently reinforces stereotypes). Awareness of these methodological biases is crucial for interpreting fairness evaluation results, as they can significantly influence conclusions about model performance across demographic groups. This comprehensive typology enables more systematic and targeted fairness evaluations. By clearly differentiating between demographic characteristic biases and dataset construction biases, researchers can better identify the specific mechanisms through which unfairness manifests in LMs. Additionally, this framework facilitates more precise dataset selection, allowing researchers to choose evaluation resources that align with their specific fairness concerns and application contexts.
2.1.2.4  Accessibility
@@
Finally, we consider accessibility, which influences how readily they can be adopted, reproduced, and extended by the research community. Datasets generally fall into two categories: (1) publicly available datasets and (2) restricted-access datasets, though the majority fall under the public category.
@@
Publicly available datasets are accessible to researchers and developers without major barriers. These datasets are often hosted on platforms like Hugging Face Datasets [^3] https://huggingface.co or GitHub repositories [^4] https://github.com and come with documentation and licensing terms that encourage open use. They are essential for promoting transparency, collaboration, and replicability in fairness research, enabling broad participation in the evaluation and improvement of LMs.
@@
2.2 Dataset Collection and Selection Criteria
2.2.1 Search Strategy
To identify relevant datasets and benchmarks, we queried four major academic databases: IEEE Xplore, ACM Digital Library, Scopus, and Google Scholar, along with the pre-print repository arXiv. Our primary search strings were “fairness datasets” and “fairness datasets”, augmented with secondary terms such as “bias”, “discrimination”, and “language models”. Because research on NLP fairness intersects AI, ethics, linguistics, and social computing, we focused on venues that consistently publish new resources and evaluations. In practice, this meant screening proceedings and journals from ACL [24, 26, 27, 52, 42, 41, 34, 35, 47, 19, 53, 54], EMNLP [25, 30, 18, 33, 45, 29, 55], NAACL [28, 51, 32, 56, 57], COLING [58, 59, 60], LREC [61], NeurIPS [62, 63, 64, 65, 66, 67], ICML [43], AAAI [68, 69, 70, 71], and ACM FAccT [36, 72]. Foundational arXiv benchmark proposals such as RealToxicityPrompts [37] and TrustGPT [43] were also examined. Finally, we performed backward‑snowballing on the reference lists of retrieved papers and consulted recent survey articles on fairness in language models [23, 16, 73, 21, 74, 75, 76, 77] to ensure comprehensive coverage.
2.2.2 Inclusion and Exclusion Criteria
Inclusion Criteria. A dataset was admitted into the corpus if it primarily targets the detection of unjustified adverse effects on specific individuals or social groups, offers a transparent and replicable account of its construction, and demonstrates practical relevance through peer citations or real-world deployments. We further required comprehensive technical documentation, including data format, label taxonomy, and annotation procedures, to ensure reproducibility, as well as an explicit connection to current research questions in LM fairness. Only datasets satisfying all of these conditions advanced to the analysis stage.
Selection Outcome. From an initial pool of 137 candidate datasets, the inclusion–exclusion filter yielded 16 datasets that satisfy all criteria for systematic fairness evaluation of LMs.
@@ 2.3 Bias Definitions and Measurement Framework
Building on the two‑level taxonomy in Section 2.1, We formalize three dataset-level bias types: representativeness bias, annotation bias, and stereotype leakage. Each is associated with a principled estimator, allowing task-agnostic quantification. These definitions and metrics form the backbone of our analysis in Sections 3 and 4, enabling consistent comparison across datasets.
To make these notions precise, we first introduce a unified notation for datasets and their protected attributes, labels, and scores. Each dataset is modelled as a finite multiset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, where an item $x_{i}$ may be a sentence, a prompt–completion pair, or any other text unit supplied to or produced by a language model. Every instance carries an optional protected attribute $a_{i}\!\in\!\mathcal{A}$ (e.g. gender, race), an optional discrete gold label $y_{i}\!\in\!\mathcal{Y}$ (classification settings), and/or a continuous reference score $s_{i}\!\in\![0,1]$ (generation‑quality or toxicity benchmarks). Let $P_{\mathcal{P}}$ denote the true population distribution over $\mathcal{A}$ that serves as the fairness baseline.
Under this notation, we define (i) representativeness bias as divergence between the empirical attribute distribution of $\mathcal{D}$ and $P_{\mathcal{P}}$, (ii) annotation bias as systematic group-conditioned differences induced by labeling or scoring functions, and (iii) stereotype leakage as identity–trait associations revealed by co-occurrence statistics. We provide their concrete estimators (e.g., $D_{\mathrm{KL}}$ for representativeness; group-wise expectation gaps for annotation; PMI/MI for stereotype leakage) immediately below and apply them uniformly to each dataset in Sections 3 and 4.
Under this notation, we define representativeness bias, annotation bias, stereotype leakage. The remainder of this subsection formalises each type and presents a principled estimator that can be computed regardless of the underlying task or label space.
@@
We define representativeness bias using the Kullback-Leibler (KL) divergence between the dataset distribution $P_{\mathcal{D}}(A)$ and population distribution $P_{\mathcal{P}}(A)$:
$B_{\mathrm{rep}}=D_{\mathrm{KL}}(P_{\mathcal{D}}(A)\;\|\;P_{\mathcal{P}}(A))=\sum_{a\in\mathcal{A}}P_{\mathcal{D}}(a)\log\frac{P_{\mathcal{D}}(a)}{P_{\mathcal{P}}(a)}$ (1)
For instance, the BBQ dataset underrepresents Indigenous identities ($2\%$) relative to U.S. census estimates ($\sim$5%), yielding a non-zero $B_{\mathrm{rep}}$.
Annotation Bias. Annotation bias arises when the labeling function, whether human or automatic, systematically assigns different labels to semantically similar inputs based on group membership. Let $g_{\theta}(x)$ be the labeling function and $A$ the protected attribute.
This bias includes both manual annotation errors (e.g., human raters exhibiting bias) and automated model bias (e.g., models like Perspective API scoring differently based on the protected attribute, even for the same content). To measure annotation bias, we estimate the maximum expected difference in label values between demographic groups:
$B_{\text{ann}}=\max_{a_{1},a_{2}\in\mathcal{A}}\left|\mathbb{E}[g_{\theta}(x)\mid A=a_{1}]-\mathbb{E}[g_{\theta}(x)\mid A=a_{2}]\right|$ (2)
@@
$\hat{B}_{\mathrm{ann}}=\max_{a_{1},a_{2}\in\mathcal{A}}\left|\mathbb{E}_{x\sim\mathcal{D}_{a_{1}}}[y]-\mathbb{E}_{x\sim\mathcal{D}_{a_{2}}}[y]\right|$ (3)
Human annotation can itself encode group-conditioned skew. For example, in the widely used hate-speech corpus of Davidson et al. [78], crowdworkers labeled tweets written in African-American English (AAE) as “offensive” substantially more often than semantically comparable tweets in Standard American English (SAE). Once such asymmetries enter a dataset, they constitute dataset-level annotation bias ($B_{\mathrm{ann}}>0$), independent of any downstream model.
Many fairness datasets lack complete gold labels or additionally rely on automated proxy raters (i.e., auxiliary scoring functions). We formalize a family of raters $\mathcal{H}=\{h_{k}\}_{k=1}^{K}$, each producing a continuous score $s_{k}(x)$ (e.g., sentiment $\in[-1,1]$, toxicity/regard $\in[0,1]$) and, optionally, a thresholded label $\tilde{y}_{k}(x)=\mathbb{I}[s_{k}(x)\geq\tau_{k}]$. For each rater we report three complementary disparities:
$\widehat{B}^{(k)}_{\mathrm{score}}\;=\;\max_{a_{1},a_{2}\in\mathcal{A}}\Big|\,\mathbb{E}_{x\sim D_{a_{1}}}[s_{k}(x)]-\mathbb{E}_{x\sim D_{a_{2}}}[s_{k}(x)]\,\Big|,$ (4)
$\widehat{B}^{(k)}_{\mathrm{rate}}\;=\;\max_{a_{1},a_{2}\in\mathcal{A}}\Big|\,\Pr_{x\sim D_{a_{1}}}\!\big[\tilde{y}_{k}(x)=1\big]-\Pr_{x\sim D_{a_{2}}}\!\big[\tilde{y}_{k}(x)=1\big]\,\Big|,$ (5)
$\widehat{B}^{(k)}_{\mathrm{dist}}\;=\;\max_{a_{1},a_{2}\in\mathcal{A}}\,W_{1}\!\big(P_{s_{k}|A=a_{1}},\,P_{s_{k}|A=a_{2}}\big),$ (6)
where $W_{1}$ is the 1-Wasserstein distance between score distributions. When minimally contrasted counterfactual pairs $(x,x^{\prime})$ exist that differ only in identity markers, we also measure the rater’s identity sensitivity,
$\widehat{B}^{(k)}_{\mathrm{cf}}\;=\;\mathbb{E}_{(x,x^{\prime})\in\mathcal{C}}\big|\,s_{k}(x)-s_{k}(x^{\prime})\,\big|,$ (7)
to separate dataset-induced disparities from rater-induced ones. For reporting, we provide the vector $\mathbf{B}_{\mathrm{ann}}^{\mathrm{proxy}}=\big[\widehat{B}^{(k)}_{\mathrm{score}},\widehat{B}^{(k)}_{\mathrm{rate}},\widehat{B}^{(k)}_{\mathrm{dist}},\widehat{B}^{(k)}_{\mathrm{cf}}\big]_{k=1}^{K}$ and its max -norm $\|\mathbf{B}_{\mathrm{ann}}^{\mathrm{proxy}}\|_{\infty}$; uncertainty is quantified via nonparametric bootstrap confidence intervals stratified by $A$. To ensure the proxy operates on coherent linguistic units, we compose a minimal utterance $u(x)$ (e.g., question + answer) before scoring.
i) Sentiment. We use VADER [78] to obtain a valence score $s_{\mathrm{sent}}(x)\in[-1,1]$; for interpretability we optionally threshold at $\geq 0.5$ (positive) and $\leq-0.5$ (negative), following prior work [36]. Sentiment is not itself a bias label; it serves as an auxiliary signal to detect group-linked polarity skews in comparable text.
ii) Toxicity. We query the Perspective API [79] to obtain calibrated probabilities for attributes such as TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, PROFANITY, and THREAT. We define a composite score $s_{\mathrm{tox}}(x)$ either as the maximum across attributes or as $1-\prod_{j}\!\big(1-s_{\mathrm{tox},j}(x)\big)$; a binary label $\tilde{y}_{\mathrm{tox}}(x)$ is produced via a threshold (e.g., $0.5$) when needed. Because Perspective is itself sensitive to domain and dialect, we always report both threshold-free distributional distances (Equation 6) and counterfactual sensitivity (Equation 7).
iii) Regard. We employ the regard classifier of Sheng et al. [80], which estimates polarity toward demographics, yielding $s_{\mathrm{reg}}(x)$ (or $\{\text{neg},\text{neu},\text{pos}\}$). As the model is trained on a limited set of groups (e.g., gender, race), we restrict analyses to supported $A$. Unlike generic sentiment, regard targets attitude toward an identity, aligning more directly with fairness concerns.
iv) Gender Polarity. We use two complementary signals following Dhamala et al. [36]. (i) Unigram matching counts gendered tokens to assign $\{\text{male},\text{female},\text{neutral}\}$, capturing explicit mentions. (ii) Embedding-based polarity leverages the gender direction $\mathbf{g}=\mathbf{she}-\mathbf{he}$ in a hard-debiased Word2Vec space [81]; for token vector $\mathbf{w}_{i}$,
$b_{i}\;=\;\frac{\langle\mathbf{w}_{i},\mathbf{g}\rangle}{\|\mathbf{w}_{i}\|\,\|\mathbf{g}\|},$ (8)
aggregated as
$\mathrm{Gender\text{-}Wavg}(x)\;=\;\frac{\sum_{i=1}^{n}\mathrm{sgn}(b_{i})\,b_{i}^{2}}{\sum_{i=1}^{n}|b_{i}|},\qquad\mathrm{Gender\text{-}Max}(x)\;=\;\mathrm{sgn}\!\big(b_{i^{\ast}}\big)\,|b_{i^{\ast}}|,\;\;i^{\ast}=\arg\max_{i}|b_{i}|.$ (9)
We convert to discrete polarity using thresholds (e.g., $\geq 0.25$ female, $\leq-0.25$ male) as in Dhamala et al. [36], while also reporting distributional distances (Equation 6) for threshold-free comparison.
@@
$\mathrm{PMI}(w_{g},w_{t})\;=\;\log\frac{P(w_{g},w_{t})}{P(w_{g})\,P(w_{t})},\qquad w_{g}\in W_{G},\;w_{t}\in W_{T}$ (10)
Corpus-level association. Mutual Information (MI) averages PMI over all group-trait pairs, yielding a single scalar that summarizes stereotype leakage for the whole dataset:
$\mathrm{MI}(W_{G};W_{T})\;=\;\sum_{w_{g}\in W_{G}}\;\sum_{w_{t}\in W_{T}}P(w_{g},w_{t})\,\log\frac{P(w_{g},w_{t})}{P(w_{g})\,P(w_{t})}$ (11)
@@
3 Constrained-Form Evaluation Datasets
Figure 2: Overview of constrained-form evaluation datasets discussed in Section 3.1.
Building on the two-level taxonomy introduced in Section 2.1, we now examine the constrained-form family of fairness datasets. In this setting, language models choose, rank, or score among predefined outputs such as classification labels, multiple-choice options, or ranked lists, which yields well-specified tasks and unambiguous measurement. The constrained format enables precise and comparable audits across demographic conditions using metrics such as accuracy gaps, probability shifts, selection rates, and rank-based disparities, and it is well-suited to diagnosing biases that arise in fixed decision processes where protected attributes should be irrelevant. To orient the reader, Figure 2 provides a visual map of the five task families and the representative datasets covered in this section; we then proceed to dataset-specific analyses organized by primary evaluation focus.